Torque can be calculated by multiplying the force (N) applied at 90∘ to the lever arm at a distance (m) from the pivot point (point of rotation), the compound SI unit for the torque is N⋅m. If the force (at 90∘ to the lever arm) applied is 15 N and it is applied at 2.0 m from the pivot point (point of rotation), what is the torque on the lever?
Answer #1

\,

The magnitude of torque is given by the equation,

\,

\tau =Fd \sin \theta

\,

inserting the data given,

\,

\tau =(15N)(2m)( \sin 90^{\circ} )

\,

or

\,

{\color{Blue} \tau =30 \, N \cdot m}